How would I solve this? I'm 46 and trying to remember algebra...
A sailboat travels 20 miles downstream in 3 hours. It returns in 4 hrs. Find the speed of the sailboat in still water and the rate of the current.

Answer by richwmiller(17219)

b=boat
c=current
3(b+c)=20
4(b-c)=20
b=35/6 c=5/6
check
3(35/6+5/6)=3*40/6=40/2=20
4*(35/6-5/6)=4*30/6=4*5=20
ok